OpenStudy (anonymous):

show how to find the zeros, and then describe the other key features of g(x) g(x) = x3 − x2 − 4x + 4

OpenStudy (anonymous):

@Mertsj

OpenStudy (campbell_st):

well grouping in parts works \[g(x) = (x^3 - x^2) -(4x - 4)\] now factor each pair... what do you get..?

OpenStudy (mertsj):

Follow up on the factoring as suggested above. Factor x^2 out of the first group. Factor 4 out of the second group

OpenStudy (welshfella):

g(x) = x^2(x - 1) - 4(x - 1) = 0

OpenStudy (welshfella):

(x^2 - 4)(x - 1) = 0 now you can factor x^2 - 4 ( difference of 2 squares) to give 3 factors which will give you 3 zeroes